# Latest version here

### v2

@@ Abstract
Keywords: signal extraction, reversion to the mean, underreaction, central tendency effect, cognitive imprecision
@@ 1 Introduction
In many economic contexts, agents fail to respond optimally to fundamental variables: firms and households adjust only partially to macroeconomic changes (Sims, 2003), consumers underreact to non-salient taxes (Chetty, Looney, and Kroft, 2009), and numerical estimates are biased toward default values (Tversky and Kahneman, 1974). [^3] Outside economics, psychophysics documents a central-tendency effect: perceptual judgments are biased toward the mean of the stimulus distribution (Hollingworth, 1910; Stevens and Greenbaum, 1966).
@@
These explanations share a core mechanism: noise attenuates Bayesian updating, which compresses behavior toward a default action. This mechanism is usually formalized using the normal-normal model, where the agent observes a noisy signal $S=X+\varepsilon$. Here, the state $X\in\mathbb{R}$ is normally distributed, and $\varepsilon\in\mathbb{R}$ is independent, normal noise. The agent’s posterior mean is then compressed toward the prior mean, and more so when the signal is less precise: the posterior mean lies between the signal and the prior mean, and the posterior mean is closer to the prior mean the larger the variance of $\varepsilon$ [^4] Formally, if ${X}\sim\mathcal{N}(\mu,\sigma_{X}^{2})$ and $\varepsilon\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2})$, then $\mathbb{E}[{X}|S=s]=\frac{\sigma_{{X}}^{2}}{\sigma_{X}^{2}+\sigma_{\varepsilon}^{2}}s+\left(1-\frac{\sigma_{{X}}^{2}}{\sigma_{{X}}^{2}+\sigma_{\varepsilon}^{2}}\right)\mu.$ For any signal realization $s$, the larger $\sigma_{\varepsilon}^{2}$, the closer the posterior mean is to the prior mean. We refer to this effect as imprecision attenuates updating. If behavior is based on the posterior mean, then such attenuation of updating leads to attenuation of behavior. In fact, a key piece of evidence for cognitive imprecision relies on this effect: subjects reporting higher cognitive uncertainty exhibit more attenuated behavioral responses (Enke and Graeber, 2023; Enke, Graeber, Oprea, and Yang, 2024).
@@
Other comparative statics results have been established for location experiments with log-concave noise distributions, which satisfy the monotone likelihood-ratio property. Milgrom (1981) shows that the strict monotone likelihood-ratio property ensures higher signals constitute “good news,” in the sense of producing first-order stochastically dominant posteriors, for any prior. [^6] See also Chambers and Healy (2011) for a strengthening of Milgrom’s result and Heinsalu (2020) for an extension and economic application of Chambers and Healy (2011). Conversely, Dawid (1973) examines violations of the “good news”-property under heavy-tailed noise distributions, showing that if the noise is sufficiently heavy-tailed relative to the prior, extreme observations are “rejected,” yielding posteriors that revert to the prior (for further discussion, see O’Hagan and Pericchi, 2012). Together with our results, this highlights log-concave location-scale experiments are a well-behaved class of signal structures that retains key properties of the normal-normal model. Kartik, Lee, and Suen (2021) provides a comparative statics result with economic applications for agents with heterogeneous priors, termed “information validates the prior,” in which a more informative experiment brings another agent’s posterior mean closer to one’s prior mean.
@@ 2 Model
We restrict attention to signal structures, where the signal equals the one-dimensional state plus some independent noise, which are called location experiments. Formally, the random signal $S$ is a location experiment if it is the sum of the random state $X$ and random noise $\varepsilon$ that is independent of ${X}$, which we write as $\varepsilon\perp\!\!\!\perp{X}$. We assume the noise $\varepsilon$ has mean 0; that is, the signal is unbiased. If not, one could easily derive an unbiased signal by subtracting the mean of $\varepsilon$ from $S$.
@@
Further, we make the technical assumption that the state ${X}$ and noise $\varepsilon$ admit positive, continuously differentiable densities $f_{X}$ and $f_{\varepsilon}$ with finite absolute first moments. This guarantees finite conditional expectation $\mathbb{E}[{X}|S=s]$. We call $f_{X}$ the prior density and $f_{\varepsilon}$ the noise density.
@@
The random state ${X}$ and noise $\varepsilon$ admit positive, continuously differentiable densities $f_{X},f_{\varepsilon}$ and have finite absolute first moments.
@@ 3 Precision Order
[…]
@@ 4.1 Location-Scale Experiments
If $\varepsilon$ is symmetric around 0 and $\log f_{\varepsilon}(e^{x})$ is concave, then $\forall\sigma^{\prime}>\sigma>0$, $\sigma^{\prime}\varepsilon$ is less precise than $\sigma\varepsilon$. In particular, $\log f_{\varepsilon}(e^{x})$ is concave if the density $f_{\varepsilon}(x)$ is log-concave. [^10] This can be seen easily if the log-density is differentiable. Define $\phi=\log f$ and $\psi=\exp$. Then, $\phi$ is concave and decreasing for positive values and $\psi$ is convex, increasing, and obtains positive values only, which implies $(\phi\circ\psi)^{\prime\prime}(x)=\psi^{\prime\prime}(x)(\phi^{\prime}\circ\psi)(x)+(\psi^{\prime}(x))^{2}(\phi^{\prime\prime}\circ\psi)(x)\leq 0$.
In the Appendix, we show that Lemma 1 follows from a known result. Many commonly used distributions are symmetric and log-concave, such as the normal, logistic, extreme value, and double-exponential distributions. Further, we give examples of symmetric and non-log-concave distributions for which $\log f_{\varepsilon}(e^{x})$ is nevertheless concave, such as the Student-t, Cauchy, and the “double” Pareto distribution.
@@ 4.2 Comparative Statics on the Prior
Let the noise density be log-concave. The posterior mean is weakly closer to the prior mean under a more precise prior, for any signal realization $s$. Formally, if $\mathbb{E}[{X}]=\mathbb{E}[\tilde{X}]$ and $\tilde{X}$ is more precise than ${X}$, then
@@
By the same argument, an analogous dual result to Corollary 1 holds for scaling the prior instead of the noise density. If ${X}$ has density $f_{X}(x)$, then $k{X}$ has density $1/kf_{X}(x/k)$.
@@ 4.3 Average Posterior Means
Our previous comparative statics results hold for any signal realization and thus speaks to situations where agents observe the same signal. However, in many situations, agents observe distinct signals from the same signal structure, that is their signal realization are independent conditional on the state. Further, often we do not observe agent’s signal realizations but only the average posterior mean, such as when we observe only aggregate behavior from a population of individuals. What can be said about comparative statics with respect to the average posterior means given some true state ${X}$?
Before we prove comparative statics results, we show that the average posterior mean necessarily lies between the state ${X}$ and the prior mean $\mathbb{E}[{X}]$, extending Fact 1.
@@
For any state ${x}$, the conditional average posterior mean $\mathbb{E}[\mathbb{E}[{X}|S]|{X}=x]$ lies weakly between the state $x$ and the prior mean $\mathbb{E}[{X}]$. Formally,
@@
The proof is illustrated using Figure 2. Let ${X}>\mathbb{E}[{X}]=0$. Conditional on state ${X}$, the distribution of the signal $S$ is is symmetric around ${X}$ so its expectation equals ${X}$. Taking the conditional expectation $\mathbb{E}[{X}|S]$ moves the distribution closer to $\mathbb{E}[{X}]$ as indicated by the arrows. Given our assumptions, the conditional expectation is antisymmetric and $s>0$ has a higher likelihood than $-s<0$. Thus, the overall effect on the average posterior mean is negative and $\mathbb{E}[\mathbb{E}[{X}|S]|{X}]<\mathbb{E}[S|{X}]={X}$. Further, because the density of $s$ is larger than the density of $-s$, integrating over all $s$ leads to a positive expectation, so $\mathbb{E}[{X}]<\mathbb{E}[\mathbb{E}[{X}|S]|{X}]$.
@@
First, we prove a comparative statics result for overconfidence in the signal. We consider two agents, $A$ and $B$, that face the same objective signal structure $S={X}+\varepsilon$ but update differently because they have different confidence in the signal. That is, agent $i\in\{A,B\}$ forms their conditional expectation $\mathbb{E}_{i}[{X}|S=s]$ as if $S={X}+\varepsilon_{i}$, $\mathbb{E}_{i}[X|S=s]:=\mathbb{E}[X|X+\varepsilon_{i}=s]$. Especially empirically relevant is the case overconfidence in the signal, also called overprecision, which is pervasive (Moore and Healy, 2008). We define a relative notion of overconfidence by generalizing the definition in Ortoleva and Snowberg (2015), which is based on the normal-normal model, using our Definition 1.
@@
$A$ is more confident than $B$ in the signal if $\varepsilon_{A}$ is more precise than $\varepsilon_{B}$.
@@
Let the prior density be log-concave and $A$ be more confident than $B$. Conditional on any state $x$, the average posterior mean of $A$ is weakly closer to the state than the average posterior mean of $B$. Formally,
$\forall x\leq\mathbb{E}[X]\colon\qquad x\leq\mathbb{E}[\mathbb{E}_{A}[{X}|S]|{X}=x]\leq\mathbb{E}[\mathbb{E}_{B}[{X}|S]|{X}=x],$
$\forall x\geq\mathbb{E}[X]\colon\qquad x\geq\mathbb{E}[\mathbb{E}_{A}[{X}|S]|{X}=x]\geq\mathbb{E}[\mathbb{E}_{B}[{X}|S]|{X}=x].$
Figure 3: Illustrating the proof of Proposition 2
This shows that more overconfident agents have, on average, posterior means closer to the state ${X}$ and further away from their prior mean $\mathbb{E}[{X}]$.
The proof builds on the proof of Proposition 1 and is illustrated using Figure 3. Because $A$ is more overconfident in the signal, their posterior mean is closer to the signal for any signal realization $s$ as well as $-s$. For positive ${X}$, because $s>0$ is more likely than $-s$, the overall effect on the average posterior mean is positive. Then, the result follows from Proposition 1.
@@
Second, we show comparative statics with respect to the prior precision. Given some state ${X}$, consider two agents, $A$ and $B$, with symmetric and quasi-concave priors that have the same mean, where agent $A$’s prior is more precise than $B$’s. Then, we have from the argument in Proposition 2 and the Corollary 2, we immediately obtain the following result.
@@
Let the noise density be log-concave and $A$’s prior be less precise than $B$’s. Conditional on any state ${x}$, the average posterior mean of $A$ is weakly closer to the state than the average posterior mean of $B$. Formally,
$\forall x\leq\mathbb{E}[X]\colon\qquad x\leq\mathbb{E}[\mathbb{E}_{A}[{X}|S]|{X}=x]\leq\mathbb{E}[\mathbb{E}_{B}[{X}|S]|{X}=x],$
$\forall x\geq\mathbb{E}[X]\colon\qquad x\geq\mathbb{E}[\mathbb{E}_{A}[{X}|S]|{X}=x]\geq\mathbb{E}[\mathbb{E}_{B}[{X}|S]|{X}=x].$

### v3

@@ Abstract
Keywords: signal extraction, reversion to the mean, underreaction, central tendency effect, anchoring and adjustment, cognitive imprecision
@@ 1 Introduction
In many economic contexts, agents fail to respond optimally to fundamental variables: firms and households adjust only partially to macroeconomic changes (Sims, 2003), consumers underreact to non-salient taxes (Chetty, Looney, and Kroft, 2009), and numerical estimates are biased toward default values (Tversky and Kahneman, 1974). [^3] Beyond economics, psychophysics documents a central-tendency effect: perceptual judgments are biased toward the mean of the stimulus distribution (Hollingworth, 1910; Stevens and Greenbaum, 1966).
@@
These explanations share a core mechanism: noise attenuates Bayesian updating and this compresses behavior toward a default action. This mechanism is usually formalized using the normal-normal model, where the agent observes a noisy signal $S=X+\varepsilon$. Here, the state $X\in\mathbb{R}$ is normally distributed, and $\varepsilon\in\mathbb{R}$ is independent, normal noise. The agent’s posterior mean is then compressed toward the prior mean, and more so when the signal is less precise: the posterior mean lies between the signal and the prior mean, and the posterior mean is closer to the prior mean the larger the variance of $\varepsilon$. [^4] Formally, if ${X}\sim\mathcal{N}(\mu,\sigma_{X}^{2})$ and $\varepsilon\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2})$, then $\mathbb{E}[{X}|S=s]=\frac{\sigma_{{X}}^{2}}{\sigma_{X}^{2}+\sigma_{\varepsilon}^{2}}s+\left(1-\frac{\sigma_{{X}}^{2}}{\sigma_{{X}}^{2}+\sigma_{\varepsilon}^{2}}\right)\mu.$ For any signal realization $s$, the larger $\sigma_{\varepsilon}^{2}$, the closer the posterior mean is to the prior mean. We refer to this effect as imprecision attenuates updating. If behavior is based on the posterior mean, then such attenuation of updating leads to attenuation of behavior. In fact, a key piece of evidence for cognitive imprecision relies on this effect: subjects reporting higher cognitive uncertainty exhibit more attenuated behavioral responses (Enke and Graeber, 2023; Enke, Graeber, Oprea, and Yang, 2024).
@@
Other comparative statics results have been established for location experiments with log-concave noise distributions, which satisfy the monotone likelihood-ratio property. Milgrom (1981) shows that the strict monotone likelihood-ratio property ensures higher signals constitute “good news,” in the sense of producing first-order stochastically dominant posteriors, for any prior. [^6] See also Chambers and Healy (2011) for a strengthening of Milgrom’s result and Heinsalu (2020) for an extension and economic application of Chambers and Healy (2011). Conversely, Dawid (1973) examines violations of the “good news”-property under heavy-tailed noise distributions, showing that if the noise is sufficiently heavy-tailed relative to the prior, extreme observations are “rejected,” yielding posteriors that revert to the prior (for further discussion, see O’Hagan and Pericchi, 2012). Together with our results, this highlights log-concave location-scale experiments are a well-behaved class of signal structures that retains key properties of the normal-normal model. Kartik et al. (2021) provides a comparative statics result with economic applications for agents with heterogeneous priors, termed “information validates the prior,” in which a more informative experiment brings another agent’s posterior mean closer to one’s prior mean.
@@ 2 Model
We restrict attention to signal structures, where the signal equals the one-dimensional state plus some independent noise, which are called location experiments. Formally, the random signal $S$ is a location experiment if it is the sum of the random state $X$ and random noise $\varepsilon$ that is independent of X{X}, which we write as $\varepsilon\perp\!\!\!\perp{X}$. We assume the noise $\varepsilon$ has mean 0; that is, the signal is unbiased. If not, one could easily derive an unbiased signal by subtracting the mean of $\varepsilon$ from $S$.
@@
Further, we make the technical assumption that the state X{X} and noise $\varepsilon$ admit positive, continuously differentiable densities $f_{X}$ and $f_{\varepsilon}$ with finite absolute first moments. This guarantees finite conditional expectation $\mathbb{E}[{X}|S=s]$. We call $f_{X}$ the prior density and $f_{\varepsilon}$ the noise density.
@@
The random state X{X} and noise $\varepsilon$ admit positive, continuously differentiable densities $f_{X},f_{\varepsilon}$ and have finite absolute first moments.
@@ 3 Precision Order
The precision order lies at the intersection of two important orders: the convex order (also called mean-preserving spread order) on noise distributions, and the Lehmann order on location experiments. One can show that if the location experiment $S=X+\varepsilon^{\prime}$ is less precise than $S=X+\varepsilon$, then noise $\varepsilon^{\prime}$ is a mean-preserving spread of noise $\varepsilon$. [^8] If $S^{\prime}$ is less precise than $S$, the CDFs of $\varepsilon^{\prime}$ and $\varepsilon$ cross once. Together with $\varepsilon^{\prime}$ and $\varepsilon$ having equal means, this implies one distribution majorizes the other; see also Diamond and Stiglitz (1974). Note that a mean-preserving spread is less demanding than the Blackwell order on location experiments, which requires that the additional noise is independent. The Lehmann order characterizes what experiments are more valuable in all monotone decision problems (Lehmann, 1988; Quah and Strulovici, 2009). This order is defined only for location experiments with log-concave noise, because only these location experiments satisfy the required monotone likelihood-ratio property. One can show for location experiments with log-concave noise, the precision order implies the Lehmann order. [^9] On such location experiments, the Lehmann order coincides with the dispersive order on the noise distributions, which requires that any two quantiles are weakly further apart under the more dispersed noise distribution. Formally, let $\varepsilon$ and $\varepsilon^{\prime}$ have CDFs $F$ and $G$, respectively, and let $F^{-1}$ and $G^{-1}$ denote here the right-continuous inverses. $\varepsilon$ is smaller in the dispersive order than $\varepsilon^{\prime}$ if $\forall\ 0<\alpha\leq\beta<1:F^{-1}(\beta)-F^{-1}(\alpha)\leq G^{-1}(\beta)-G^{-1}(\alpha).$ When we truncate the noise to positive values, the precision order implies the likelihood-ratio order. As is known, the likelihood-ratio order implies the hazard-rate order (e.g., Shaked and Shanthikumar, 2007, Theorem 1.C.1). The hazard-rate order together with densities being log-concave implies the dispersive order (Bagai and Kochar, 1986, Theorem 1). One can easily see that if the dispersive order holds for all $1/2\leq\alpha\leq\beta<1$, then it holds for all $0<\alpha\leq\beta<1$.
@@ 4.1 Location-Scale Experiments
If $\varepsilon$ is symmetric around 0 and $\log\varepsilon$ has a log-concave density, then $\forall\sigma^{\prime}>\sigma>0$, $\sigma^{\prime}\varepsilon$ is less precise than $\sigma\varepsilon$.
It is worth noting that $\log\varepsilon$ has a log-concave density if $\varepsilon$ has a log-concave density. [^10] This can be seen easily if the log-density is differentiable. Define $\phi=\log f$ and $\psi=\exp$. Then, $\phi$ is concave and decreasing for positive values and $\psi$ is convex, increasing, and obtains positive values only, which implies $(\phi\circ\psi)^{\prime\prime}(x)=\psi^{\prime\prime}(x)(\phi^{\prime}\circ\psi)(x)+(\psi^{\prime}(x))^{2}(\phi^{\prime\prime}\circ\psi)(x)\leq 0$. In the Appendix, we show that Lemma 1 follows from a known result. Many commonly used distributions are symmetric and log-concave, such as the normal, logistic, extreme value, and double-exponential distributions. Further, we give examples of symmetric and non-log-concave distributions for $\varepsilon$ for which $\log\varepsilon$ is log-concave, such as the Student-t, Cauchy, and the “double” Pareto distribution.
@@ 4.2 Comparative Statics on the Prior
Let the noise density be log-concave. The posterior mean is weakly closer to the prior mean under a more precise prior, for any signal realization $s$. Formally, if $\mathbb{E}[{X}]=\mathbb{E}[\tilde{X}$] and $\tilde{X}$ is more precise than X{X}, then
@@
By the same argument, an analogous dual result to Corollary 1 holds for scaling the prior instead of the noise density. If X{X} has density $f_{X}(x)$, then kXk{X} has density $1/kf_{X}(x/k)$.
@@ 4.3 Average Posterior Means
Our previous comparative statics results hold for any signal realization and thus speaks to situations where agents observe the same signal. However, in many situations, agents observe distinct signals from the same signal structure, that is their signal realization are independent conditional on the state. Further, often we do not observe agent’s signal realizations but only the average posterior mean, such as when we observe only aggregate behavior from a population of individuals. What can be said about comparative statics with respect to the average posterior means given some true state?
Before we prove comparative statics results, we show that the average posterior mean necessarily lies between the state X{X} and the prior mean $\mathbb{E}[{X}]$, extending Fact 1.
@@
For any state x{x}, the conditional average posterior mean $\mathbb{E}[\mathbb{E}[{X}|S]|{X}=x]$ lies weakly between the state $x$ and the prior mean $\mathbb{E}[{X}]$. Formally,
@@
The proof is illustrated using Figure 2. Let the state be greater than expected and normalize the expectation to 0, $x>\mathbb{E}[{X}]=0$. Conditional on $X=x$, the distribution of the signal $S$ is is symmetric around $x$ so its expectation equals $x$. Pushing forward the conditional distribution of the signal through the conditional-expectation function $\mathbb{E}[{X}|S]$ moves the distribution closer to $\mathbb{E}[{X}]=0$ as indicated by the arrows. Given our assumptions, the conditional expectation is antisymmetric and $s>0$ has a higher likelihood than s<0-s<0. Thus, the overall effect on the average posterior mean is negative and $\mathbb{E}[\mathbb{E}[{X}|S]|{X}=x]<\mathbb{E}[S|{X}]=x$. Further, because the density of $s$ is larger than the density of s-s, integrating over all $s$ leads to a positive expectation, so $\mathbb{E}[{X}]<\mathbb{E}[\mathbb{E}[{X}|S]|{X}=x]$.
@@
First, we prove a comparative statics result for overconfidence in the signal. We consider two agents, $A$ and $B$, that face the same objective signal structure $S={X}+\varepsilon$ but update differently because they have different confidence in the signal. That is, agent $i\in\{A,B\}$ forms their conditional expectation $\mathbb{E}_{i}[{X}|S=s]$ as if $S={X}+\varepsilon_{i}$, $\mathbb{E}_{i}[X|S=s]:=\mathbb{E}[X|X+\varepsilon_{i}=s]$. Especially empirically relevant is the case overconfidence in the signal, also called overprecision, which is pervasive (Moore and Healy, 2008). We define a relative notion of (over)confidence by generalizing the definition in Ortoleva and Snowberg (2015), which is based on the normal-normal model, using our Definition 1.
@@
AA is more confident than $B$ in the signal if $\varepsilon_{A}$ is more precise than $\varepsilon_{B}$.
@@
Let the prior density be log-concave and $A$ be more confident in the signal than $B$. Conditional on any state $x$, the average posterior mean of $A$ is weakly closer to the state than the average posterior mean of $B$. Formally,
$\forall x\leq\mathbb{E}[X]\colon\qquad x\leq\mathbb{E}[\mathbb{E}_{A}[{X}|S]|{X}=x]\leq\mathbb{E}[\mathbb{E}_{B}[{X}|S]|{X}=x]\leq\mathbb{E}[X],$
$\forall x\geq\mathbb{E}[X]\colon\qquad x\geq\mathbb{E}[\mathbb{E}_{A}[{X}|S]|{X}=x]\geq\mathbb{E}[\mathbb{E}_{B}[{X}|S]|{X}=x]\geq\mathbb{E}[X].$
Figure 3: Proof of Proposition 2
The result states that the average posterior mean of the more confident agent is closer to the true state X{X} and further away from the prior mean $\mathbb{E}[{X}]$.
We sketch the proof using Figure 3. We suppose again that $x>\mathbb{E}[X]=0$. Because $A$ is more confident in the signal, their posterior mean is closer to the signal for signal realization $s$ as well as s-s. Thus, the posterior mean of $A$ is sometimes greater and sometimes less than the posterior mean of $B$. However, for $x>0$, because $s>0$ is more likely than s-s, the positive effect dominates and the average posterior mean of $A$ is greater than the one of $B$.
@@
Second, we show comparative statics with respect to the prior precision. Given some state X{X}, consider two agents, $A$ and $B$, with symmetric and quasi-concave priors that have the same mean, where agent $A$’s prior is more precise than $B$’s. Then, using the proof of Proposition 2 and Corollary 2, we obtain the following result.
@@
Let the noise density be log-concave and $A$’s prior be less precise than $B$’s. Conditional on any state x{x}, the average posterior mean of $A$ is weakly closer to the state than the average posterior mean of $B$. Formally,
$\forall x\leq\mathbb{E}[X]\colon\qquad x\leq\mathbb{E}[\mathbb{E}_{A}[{X}|S]|{X}=x]\leq\mathbb{E}[\mathbb{E}_{B}[{X}|S]|{X}=x]\leq\mathbb{E}[X],$
$\forall x\geq\mathbb{E}[X]\colon\qquad x\geq\mathbb{E}[\mathbb{E}_{A}[{X}|S]|{X}=x]\geq\mathbb{E}[\mathbb{E}_{B}[{X}|S]|{X}=x]\geq\mathbb{E}[X].$
Prior precision has the opposite effect from confidence in the signal. A more precise prior leads to an average posterior mean closer to the prior mean and further from the signal.